OpenStudy (anonymous):

Find the derivative of the fcn: \[\sqrt{2x + \sqrt{2x + \sqrt{2x}} }\] Attempt: Is this the same as: \[(2x + (2x + (2x)^{1/2})^{1/2})^{1/2}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, you are going to use the chain rule, three times, in order to find this derivative.

OpenStudy (anonymous):

Ok Let me show you what I got. One minute, please get back to me soon.

OpenStudy (anonymous):

Ok here it is:

OpenStudy (anonymous):

\[f'(x) = 1/2(2x(2x + (2x)^{1/2})^{1/2})^{-1/2} * x(2x + (2x)^{1/2})^{-1/2}*(2 + x^{-1/2}) \]

OpenStudy (anonymous):

Here is what I got: \[f \prime(x)=\left[ 2x+(2x+(2x)^{1/2} \right]^{-1/2}\left[ 2+1/2(2x+(2x)^{1/2})^{-1/2} \right]\left[ 2+(2x)^{-1/2} \right]\]

OpenStudy (anonymous):

That was lengthy, and labrynthine!